My question is: A rectangular soccer field is twice as long as it is wide. If the perimeter of the soccer field is 300 yards , what are its dimensions?
I know the basic formula is 2W+2L=300 but i am not sure where to go from there...

Answer by stanbon(75887)

A rectangular soccer field is twice as long as it is wide. If the perimeter of the soccer field is 300 yards , what are its dimensions?
I know the basic formula is 2W+2L=300 but i am not sure where to go from there...
-----
Equations:
2W + 2L = 300
L = 2W
----
Substitute for "L" and solve for "W":
2W + 2(2W) = 300
6W = 300
W = 50 yds (width)
----
Solve for "L":
L = 2W
L = 100 yds (length)
========================
Cheers,
Stan H.